OpenStudy (anonymous):

The value of 12−−√ lies between which two consecutive integers?

OpenStudy (accessdenied):

You mean \( \sqrt{12} \) ? Well for any integer you have: \( n = \sqrt{n^2} \) So this question equates to finding the nearest perfect square below 12 and above 12. Those are numbers which have integer square roots. Does that make sense?

OpenStudy (accessdenied):

\( \sqrt{a^2} < \sqrt{12} < \sqrt{b^2} \) <-- a^2 < 12 < b^2 is equivalent to \( a < \sqrt{12} < b \) <-- a, b are the integers we want.